How much money do you need to invest at 2.8% compound continuously in order to have 25500 at the end of 8 years?
Answer by stanbon(75887)

How much money do you need to invest at 2.8% compound continuously in order to have 25500 at the end of 8 years?
-----
A(t) = P*e^(r*t)
----------
25500 = P*e^(0.028*8)
P = 25500/e^(2.24)
P = 25500/9.3933
P = $2714.70
--------------
Cheers,
Stan H.
----------